OpenStudy (anonymous):

Hillary has the following scores on her math tests: 65, 86, 95, 78. How much would her mean change if the lowest test score is dropped? (round to the nearest whole number) 5 points 4 points 3 points 2 points

OpenStudy (anonymous):

Th mean of the four scores is\[(65+86+95+78)\div4=324\div4=81\]By dropping the low score the average of the highest three is \[(324-65)\div3=259/3=86.333\]Thus the mean would change 86.333-81=5.333; to the nearest integer, 5 points.